A valve in a full 6000 gallon water tank is slowly opening. Water flows out of the tank through the valve. The flow rate in gallons per hour is given by the function f(t)=150t^2 where t is in minutes.
a) How much water flows out of the tank in the first 3 minutes?
b)How many minutes does it take for the tank to empty completely?

I am completely lost any help would be great!!!!

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

A valve in a full 6000 gallon water tank is slowly opening. Water flows out of the tank through the valve. The flow rate in gallons per hour is given by the function f(t)=150t^2 where t is in minutes.
a) How much water flows out of the tank in the first 3 minutes?
b)How many minutes does it take for the tank to empty completely?
---------------
What class are you in?
Can you use derivatives and integrals? Are you familiar with those?

Answer by richwmiller(17219)

a)
f(3)=150*3^2
f(3)=150*3*3=450*3=1350
f(3)=1350
b)
6000=150*t^2
40=t^2
2sqrt(10)=t
t=6.32456 minutes